a. (9-5)-(6-8)=
b. Which expression has the smallest value (-4), (-8), (7) or (16)?
Thank you so much!

Answer by richard1234(7193)

a. Do the operations in parentheses first...e.g. 9-5 = 4, 6-8 = -2. Then (9-5)-(6-8) = 4-(-2) = 4+2 = 6.

b. -8 is the smallest of the four numbers.